Grade 12th passMechanics

When the speed of a moving object is doubled its KE increased to 4 times how plz explain

Rafnas Pm
9 Years agoGrade 12th pass

1 Answer

Kenny Toijam
9 Years ago
Just apply KE=1/2 mv^2............... putting v=2v......the new KE becomes KE=1/2m(2v)^2= 1/2m*4v^2= 4KE